OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7).

OpenStudy (anonymous):

slope=y2-y1/x2-x1

OpenStudy (anonymous):

y=1/3(13-2x y=-2x/3+13/3

OpenStudy (anonymous):

What do you mean which? I can write an equation in slope intercept from for you if that's what you mean. Here is the equation: First we need to find the slope, to do this there are multiple different ways but I'm going to show you the easiest one in my opinion. To find the slope put the to points on top of one another, like this (-4, 7) (5, 1) then subtract the points (remember, rise over run, which means y goes over x, like this y/x) 6/-9 <- This is the slope. The Y-intercept is 3. now we just need to plug the numbers into the equation. slope intercept form is y = mx + b y = 6/-9x + 3

OpenStudy (anonymous):

acn you help me with one more!!

OpenStudy (anonymous):

*can

OpenStudy (anonymous):

Sure, what do you need :)